OpenStudy (anonymous):

help! h(x)sqrtrt(3x+5) for sqrth(1)

OpenStudy (anonymous):

h(x)=\[\sqrt{3x+5}\] for \[for \sqrt{h(1)}\]

hartnn (hartnn):

and what did u try for this ?

hartnn (hartnn):

could u find h(1) ?

OpenStudy (anonymous):

\[\sqrt{\sqrt{3}+5}\] i did something like this and i don't know what to do further!

hartnn (hartnn):

to find f(1), just put x=1 in h(x)

hartnn (hartnn):

***to find h(1)

OpenStudy (anonymous):

then that would be \[\sqrt{8}\]

hartnn (hartnn):

yes, and sqrt h(1) = sqrt (sqrt 8)

hartnn (hartnn):

at the most you can simplify it to sqrt (2*sqrt 2) or keep the final answer as sqrt {sqrt 8} only...

OpenStudy (anonymous):

oh ok thanks again :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

:)